OpenStudy (xmoses1):

The equations of two lines are x - 3y = 6 and y = 3x + 2. Determine if the lines are parallel, perpendicular or neither. parallel perpendicular neither

OpenStudy (xmoses1):

@Xmoses1 I need your help!

OpenStudy (xmoses1):

Yes how can I help you?

OpenStudy (xmoses1):

Funny... You are the only one that seems willing to help. There is something wrong with this picture though...

OpenStudy (xmoses1):

Well i sure wish that i had somebody other than myself to help me!

OpenStudy (xmoses1):

this blows...

OpenStudy (anonymous):

do you have a graphing calculator

OpenStudy (anonymous):

nevermind i don't think it is necessary

OpenStudy (loser66):

lol I like the guys names Xmoses1

OpenStudy (xmoses1):

no i dont do i need one?

OpenStudy (jdoe0001):

solve both for "y" firstly, what do you get for each?

OpenStudy (anonymous):

y = 3x + 2 (1) x - 3y = 6 -3y = 6 - x y = x/3 + 2 (2) Parallel = same slope, Perpendicular = negative recip. It's neither of those, so neither.

OpenStudy (anonymous):

When it is Parallel, the slopes are the same, and the y-intercept can be different. When it is perpendiculator, the slopes are a negative reciprocal if it is neither,the slopes are different

OpenStudy (xmoses1):

Ok thanks guys! Next time I will ask you guys for help rather than @Xmoses1. He was useless!